OpenStudy (pokemon23):

Saifoo I NEED YOUR HELP IN MATH ;) Saifoo I NEED YOUR HELP IN MATH ;) @Mathematics

OpenStudy (pokemon23):

Yes

OpenStudy (saifoo.khan):

yay! =D

OpenStudy (anonymous):

Saifoo!

OpenStudy (saifoo.khan):

anan!

OpenStudy (pokemon23):

express in simplest form the perimeter of a rectangle whose length is represented by 2u+3 and whose width is represented by u-1

OpenStudy (saifoo.khan):

2( 2u+3+w-1)

OpenStudy (pokemon23):

I love free medals XD !!!!!!!!!!!!!

OpenStudy (saifoo.khan):

\[4 u+2 w+4 \]

OpenStudy (pokemon23):

second question

OpenStudy (pokemon23):

Express in simplest form the area of a trapezoid whose bases are 7 and 5 and whose height is represented is represented by 8n.

OpenStudy (precal):

A=(1/2)h(b1+b2) A=(1/2)8n(7+5) A=6(8n) A=48n